### v1

# Decode Neural signal as Speech

Abstract
Decoding language from brain dynamics is an important open direction in the realm of brain-computer interface (BCI), especially considering the rapid growth of large language models. Compared to invasive-based signals which require electrode implantation surgery, non-invasive neural signals (e.g. EEG, MEG) have attracted increasing attention considering their safety and generality. However, the exploration is not adequate in three aspects: 1) previous methods mainly focus on EEG but none of the previous works address this problem on MEG with better signal quality; 2) prior works have predominantly used “teacher-forcing” during generative decoding, which is impractical; 3) prior works are mostly “BART-based” not fully auto-regressive, which performs better in other sequence tasks. In this paper, we explore the brain-to-text translation of MEG signals in a speech-decoding formation. Here we are the first to investigate a cross-attention-based “whisper” model for generating text directly from MEG signals without teacher forcing. Our model achieves impressive BLEU-1 scores of 60.30 and 52.89 without pretraining & teacher-forcing on two major datasets (GWilliams and Schoffelen). This paper conducts a comprehensive review to understand how speech decoding formation performs on the neural decoding tasks, including pretraining initialization, training & evaluation set splitting, augmentation, and scaling law.
1 Introduction
Decoding language from brain activity is an important and rapidly advancing area of neurotechnology. The ability to infer a person’s intended words, phrases, or even full sentences directly from their brain signals has profound applications for improving communication and control capabilities for those suffering from severe speech and motor impairments. Conditions like high-level spinal cord injury or late-stage ALS can trap a person inside their own body, unable to fully express themselves or interact with the digital world around them. By noninvasively measuring and modeling the complex neural representations underlying language production in the brain, we aim to restore digital communication simply through thought. Beyond assisting those with disabilities, proficient decoding of full linguistic expressions from brain signals could also enable entirely new modes of seamless human-machine interaction. One could silently yet precisely command prosthetic devices, control software applications, browse the internet, or navigate virtual and augmented reality environments entirely using their brain. The potential applications seem nearly limitless and could vastly expand what is possible for both able-bodied individuals and those with severe impairments. It is an enormously challenging task from an artificial intelligence and neuroscience perspective but one with immense promise to transform lives if achieved.
Notably, using decoding speech on invasive signals achieved high-performance recognition accuracy for real-time translation (Metzger et al., 2023; Willett et al., 2023). However, the invasive recordings pose medical risks and issues with maintaining over long periods. Thus, for non-invasive brain recordings, Wang et al. (Wang & Ji, 2022) demonstrated open-vocabulary EEG-to-text translation by employing pre-trained language models on word-level EEG features. Duan et al. (Duan et al., 2023) improved this approach by decoding from raw EEG waves without using time markers. However, limitations remain as evaluation relies on teacher forcing and the model cannot generate meaningful sentences without it. On MEG, previous works can merely decode limited classes from MEG signal. (Dash et al., 2020; Csaky et al., 2023; Ghazaryan et al., 2023). Recently, Defossez et al.(Défossez et al., 2023) decoded speech perception of 3 second segments from MEG signals, and was capable of classifying sentences by matching MEG signal with speech. However, it cannot generate sentences from the MEG signal.
To address the issue of teacher forcing evaluation and generating text from MEG signal. In this paper, we present NeuSpeech, a pioneering framework that decodes neural as speech. NeuSpeech uses an encoder-decoder model (whisper (Radford et al., 2023)) rather than BART-based models that are widely used in previous works (Wang & Ji, 2022; Duan et al., 2023; Xi et al., 2023) to directly translate raw MEG waves into text, without pretraining or transforming neural signals to discrete codes. This approach 1) lets the model efficiently learn text-related information from neural signals, and 2) keeps the general meaning of the whole sentence when generating long texts.
Experiments employ non-invasive MEG signals from the large-scale public GWilliams (Gwilliams et al., 2023) and Schoffelen (Schoffelen et al., 2019) datasets, which recorded MEG during speech listening tasks. Notably, NeuSpeech can generalize to all languages and equipment types. Performance is assessed using standard translation metrics (Papineni et al., 2002; Lin, 2004). For raw MEG waves without event markers, NeuSpeech achieves 60.30 BLEU-1 and 55.26 Rouge-1 F-measure on GWilliams, and 53.16 BLEU-1 on Schoffelen without teacher forcing evaluation. The main contributions of this work are as follows:
  • NeuSpeech adoptes an end-to-end neural framework for open-vocabulary MEG-to-text translation without markers without pre-training or teacher forcing, and is the first to achieve high-performance raw MEG wave-to-text translation.
  • We explore multi-layout, multi-language, multi-subject MEG-to-text transcription on multiple datasets under various training settings. This provides a more comprehensive evaluation than prior work.
  • We exploit cross-attention to get global semantic information for sentences rather than previous token-based methods.
2 Related Works
Significant progress has been made in the field of brain-to-text decoding in recent years. In 2019, Anumanchipalli et al. (Anumanchipalli et al., 2019) proposed a model that could decode ECoG signals into the kinematics of articulation, followed by acoustic features (such as MFCCs), and ultimately generate the speech waveform, which inspired many researchers to explore this area further. In 2020, Wang et al. (Wang et al., 2020) introduced the use of generative adversarial networks (GANs) to decode ECoG signals into speech. In 2021, Willett et al. (Willett et al., 2021) proposed a model that utilized an RNN and a probabilistic language model to decode letters while a subject was performing handwriting. Metzger et al. (Metzger et al., 2022) developed a pipeline that decoded ECoG signals into text using an RNN and then refined the results using the GPT-2 model. In the domain of open-vocabulary decoding, Metzger et al. (Metzger et al., 2023) proposed an RNN framework that could decode speech, text, emotion, and facial expressions from ECoG signals in real-time. Simultaneously, Willett et al. (Willett et al., 2023) successfully decoded text from brain signals. Liu et al. (Liu et al., 2023) developed a three-part model to decode logosyllabic languages (e.g., Chinese) by first decoding ECoG signals into Chinese pinyin with tones and syllables, and then synthesizing speech. Feng et al. (Feng et al., 2023) also achieved text decoding from SEEG signals. However, it is important to note that these practical systems predominantly rely on invasive brain recordings.
For non-invasive recordings, Meta proposed a brain-to-speech framework using contrastive learning with MEG and EEG signals (Défossez et al., 2023), which is capable of classifying limited sentences but cannot be used for open-vocabulary text decoding. Ghazaryan et al. (Ghazaryan et al., 2023) attempted to decode limited words using MEG responses. Wang et al. (Wang & Ji, 2022) developed a framework to decode word-level EEG signal features into text using a pre-trained BART model (Lewis et al., 2020). Many subsequent works have followed in the footsteps of Wang et al. (Wang & Ji, 2022), including Dewave (Duan et al., 2023), which proposed a framework that employed wave2vec (Baevski et al., 2020), and discrete codex to obtain effective representations and fed them into a BART (Lewis et al., 2020) model to generate text. However, these methods rely on teacher forcing, which is not feasible in real-world settings. Different from these methods, our work uses an end-to-end paradigm to generate open-vocabulary text and evaluate it without teacher forcing.
Table 1: Evaluation metrics of MEG-to-Text translation under raw waves input. For a fair comparison, these results has w/ and w/o the teacher-forcing evaluation setting since EEG-to-Text (Wang & Ji, 2022) only has teacher-forcing originally. Default is w/o teacher forcing and ‘w/ tf’ means with teacher-forcing.
BLEU-N (%) normal-$\uparrow$ ROUGE-1 (%)normal-$\uparrow$ WER (%) normal-$\uparrow$
Modality Method N=1 N=2 N=3 N=4 F P R
Speech Whisper (Radford et al., 2023) 67.79 59.15 51.97 45.78 75.90 73.53 81.24 41.21
Noise NeuSpeech 0.0 0.0 0.0 0.0 2.84 23.9 1.5 99.98
- random 4.99 1.18 0.59 0.41 4.79 5.95 5.87 132.30
MEG eeg-to-text (Wang & Ji, 2022) 9.21 2.13 0.57 0.14 9.74 10.73 11.38 118.25
MEG NeuSpeech 57.68 52.76 48.84 45.42 60.24 61.86 61.92 58.89
MEG NeuSpeech w/ pp 60.3 55.26 51.24 47.78 58.73 60.88 59.76 56.63
MEG eeg-to-text (Wang & Ji, 2022) w/ tf 21.66 9.27 3.93 1.73 20.11 22.48 18.71 95.69
MEG NeuSpeech w/ tf 78.13 71.25 64.79 58.86 81.91 80.52 83.68 25.38
MEG NeuSpeech w/ pp tf 80.20 73.27 66.79 60.81 78.62 78.14 79.54 23.42
3 Method
NeuSpeech is a modification of whisper (Radford et al., 2023) model, where the raw MEG signals are sent into the model and generate text. Whisper (Radford et al., 2023) is the most powerful and noise-robust model in speech recognition which is trained on hundreds of thousands of audio data and can transcribe speech from most of the languages on earth.
3.1 Task Definition
Given a raw MEG waves $\mathcal{X}$, the aim is to decode the corresponding open-vocabulary text tokens $\mathcal{W}$. These MEG-Text pairs $\langle\mathcal{X},\mathcal{W}\rangle$ are collected during hearing, as defined in Section 4.1. We decode neural signals as speech, and we only have one setting which is: Raw MEG Waves to Text Translation, where MEG waves $\mathcal{X}$ are directly sent into the whisper model to generate continuous words without teacher forcing, a more challenging but practical setting. NeuSpeech is the pioneering work in this task. Different from previous works(Duan et al., 2023; Wang & Ji, 2022), we are the first to evaluate generated text without teacher forcing using MEG.
3.2 Preprocessing
We only implement basic pre-processing here, the MEG signals are notched at power line frequency and filtered with (1,60) Hz. Then, both of them are resampled to 200Hz. After that, the signals are normalized using robust scaler mentioned in (Défossez et al., 2023), then clipped to ±$\pm 10$ and scaled into a value range of $(-1,1)$.
3.3 Model
Different from previous works (Duan et al., 2023; Tang et al., 2022) which are usually trained through multi-steps. Our model is trained end-to-end from scratch without bells and whistles. We freeze parameters in the decoder and only train the encoder. Since the original input channel of the whisper model is incompatible with the MEG channels, we replace the first convolution layer of the model with two convolution layers to fit in MEG channels and align time width respectively. For training data, we do not use any data augmentation. In terms of efficiency, we use Adalora (Zhang et al., 2023) technique to tune our model, and only the encoder including convolution layers are in the Lora model.
3.4 Evaluation
We evaluate transcribing performance on GWilliams dataset (Gwilliams et al., 2023) and Schoffelen (Schoffelen et al., 2019) using NLP metrics, BLEU, ROUGE and WER, which are used in previous works (Wang & Ji, 2022; Duan et al., 2023). We add three baselines to give a rough understanding of up and down limits. 1) speech recognition, we use a pre-trained whisper base model to demonstrate the high performance on speech modality. 2) input noise, we input Gaussian noise into our model to check if our model can distinguish between notoriously noisy MEG and noise, and the well-trained model should not generate text with pure noise as input. 3) random selection, we also showed the performance of randomly selecting sentences from training sets as predictions to give the down limit of the dataset as learned models should pass this limit to prove its effectiveness.
4 Experiments
4.1 Dataset
We utilize GWilliams (Gwilliams et al., 2023) and Schoffelen (Schoffelen et al., 2019) datasets which are auditory MEG task for experiments. Both have a detailed timetable for each sentence, and they are divided by sentence-recording pairs into train:validation:test 8:1:1, which results in 23339:2917:2918 and 8570:1112:1076 MEG-text pairs respectively. However, we should note that the unique sentences are 661 and 360 for two datasets, which are very limited compared to speech datasets.
Table 2: Transcription results on the unseen MEG waves, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings. For a fair comparison, these results do not keep the teacher-forcing evaluation setting as EEG-to-Text (Wang & Ji, 2022) and Dewave (Duan et al., 2023). This means the decoding process can hold on accumulated errors and predicts the next token with model-generated tokens from previous steps.
Decoding Results on GWilliams (Gwilliams et al., 2023)
Ground Truth: At best they tales at worst expeditions into macabre realms no healthy mind need ever see
Prediction: At best it is macred at worst expeditions into macabre realms no healthy mind need ever see that
Ground Truth: He turned to regard the pillar still rising behind his house and ran to see what had been done
Prediction: He-turned to regard a pillar still rising behind his house and ran away from…to see what had been done.
Decoding Results on Schoffelen (Schoffelen et al., 2019)
Ground Truth: Tijdens de rumoerige bespreking beslisten de leden de staking voort te zetten
Prediction: Tijdens de rumoerige bespreking beslisten de leden of de staking voort.
Ground Truth: Op die goede prestatie was helemaal niets af te dingen
Prediction: Op die goede prestatie was naar een varken afkomen van het nieuwe afguk…
4.2 Implementation Details
Model modification. We adopt whisper base model (Radford et al., 2023) as our base model. To adapt the input size of the neural signal, we use two convolution1d layers, the kernel sizes are 3, and stride is 1,2 respectively to replace the first convolution layer in the whisper-base model of dimension 512.
Training. To implement eeg-to-text (Wang & Ji, 2022), we use the original word timestamps provided by GWilliams (Gwilliams et al., 2023) to segment MEG waves for each word feature. The training settings are the same as the origin code.
All models are trained on Nvidia H800 GPUs. We use a learning rate of 1e-3 and batch size of 64 for training 500 epochs and select the best model on evaluation loss. We use the AdamW as the optimizer for training all the models.
Evaluation. For non-teacher forcing generation configuration, we use the beam search with a beam size of 5 and set the repetition penalty to 5, with no repeat n-gram size to 2, which is very important to avoid repeating. In addition, we implement post-processing composed of filtering and lowering alphabets to improve performance.
4.3 Metrics
As shown in Table 1, we compare our results to input Gaussian noise to the NeuSpeech model and randomly select sentences from the test set and whisper base model speech transcription to give the understanding of up and down limits. We also compare previous method eeg-to-text (Wang & Ji, 2022).
Based on the information provided in Tab. 1, it is evident that our method outperforms randomly selected sentences, noise input, and eeg-to-text (Wang & Ji, 2022). NeuSpeech has demonstrated the feasibility of decoding text from non-invasive brain recordings. Prior to NeuSpeech, many previous works (Duan et al., 2023; Wang & Ji, 2022) in this field relied on teacher forcing to evaluate their models, Defossez et al. (Défossez et al., 2023) used classification accuracy. However, NeuSpeech introduces a more fair and real-world setting by employing end-to-end training, which directly ties neural signals to text. In comparison to previous works evaluated using teacher forcing, NeuSpeech has shown remarkable improvement, achieving BLEU-1, BLEU-2, BLEU-3, and BLEU-4 scores of 60.3, 55.26, 51.24, and 47.78, respectively, which is close to the performance of speech recognition.
To ensure a fair comparison, we also implement teacher forcing for both the eeg-to-text model (Wang & Ji, 2022) and NeuSpeech. It is evident that our BLEU-1 score of 80.20 is almost four times that of the eeg-to-text (Wang & Ji, 2022) model, highlighting the superiority of our approach. Furthermore, we have surpassed speech recognition methods with teacher forcing, emphasizing that our model can significantly improve performance by eliminating accumulated errors. It is worth noting that, at this stage, BLEU is the most suitable metric for evaluating the MEG-to-text system, as ROUGE and WER tend to yield high scores for noise and random models.
4.4 Generated Samples
Table 3: Transcription results comparison with different methods. Bold font means exact match of words. Underline for similar but not exact matching. w/ means with. tf means teacher-forcing, pt means pretraining, jt means joint-training. eeg-to-text (Wang & Ji, 2022) here is using beam search as the same setting of ours which is described in Sec. 4.2.
(1) GT
Arthur replaced the tray onto the lacquered cherry table that stood in the center of his ordered chamber
NeuSpeech
The rental car records pointed to lacquered cherry-schedung and stood in the center of his ordered chamber in motel from Acres comb in?
NeuSpeech w/ tf
Arthur replaced thethe cartray onto thethe lacquered cherry-table inthat stood in thethe center of his ordered chamber in
NeuSpeech w/ pt
Arthur replaced the tray onto the lacquered cherry table to match from Acres.
NeuSpeech w/ jt
The rental car records pointed to Acres and stay in the center of his ordered chamber
eeg-to-text (Wang & Ji, 2022)
He said he had no idea how long it would take him to get there
eeg-to-text (Wang & Ji, 2022) w/ tf
was the old with the tablekeyered table table and had in the middle of the room table
(2) GT
My patience was long gone and I was back in the car to warming up when Acres tapped on the window and told me he had found whatever he was looking for
NeuSpeech
My patience was long gone and I was back in the car. But when I heard that many of you were looking for whatever it was, but what about this?
NeuSpeech w/ tf
My patience was long gone and I was back in athe car.to warming up when Acres wastapped on thethe window and openedtold me he had found whatever he was looking for.
NeuSpeech w/ pt
My patience was long gone and I was back in the car.
NeuSpeech w/ jt
My patience was long gone and I was back in the car. The window was looking good when Acres is ongoing on her car and she now has…
eeg-to-text (Wang & Ji, 2022)
He said he had no idea how long it would take him to get back home
eeg-to-text (Wang & Ji, 2022) w/ tf
father was running gone I was ready to the dark drive up Iacia me the window said me to was a a he was looking for
(3) GT
Very well said Arthur then take your story to that and let him defend you when no civilized publisher will approach you
NeuSpeech
Very well said Arthur! Tell me your story, and I’ll defend it from you when no civilized publisher will approach you.
NeuSpeech w/ tf
Very well said Arthur!then thattake your story to andthat and let him defend you and no civilized publisher will approach you when
NeuSpeech w/ pt
Very well said Arthur then take your story to that and let him defend you when no civilized publisher will approach you very well
NeuSpeech w/ jt
Very well said! It was that and we can make sure that it and not let him defend you when no civilized publisher will approach you
eeg-to-text (Wang & Ji, 2022)
It was the first time I had seen him in a long time and it made me very nervous
eeg-to-text (Wang & Ji, 2022) w/ tf
little then he the seat and the place go it take it he one man would you
In Tab. 2, we present visual examples of text that have been generated from MEG signals. The training of Schoffelen (Schoffelen et al., 2019) is described in sec. 4.5.1 These examples showcase the capabilities of our model in generating meaningful text, despite the difficulties associated with translating thoughts and the limited prior research in this area. While our model may not yet achieve the same level of performance as nowadays speech recognition tasks, it demonstrates the ability to align keywords and construct sentences with similar structures. This is a promising indication of the potential of our approach in the field of brain-to-text translation.
The model’s capability to match long sequences of words continuously is indeed impressive. As evidenced by the examples provided in Tab. 2, such as the first sentence with 12 identical words, the model demonstrates the ability to accurately predict groups of words throughout a sentence. It is noteworthy that the model can leverage information from the MEG signals to make predictions, allowing it to generate correct groups of words even when there are occasional errors in the middle of a sentence. This highlights the model’s capacity to capture and utilize contextual information from the brain signals, enabling it to generate meaningful and coherent text. Also, the model is good at transcribing Dutch which also shows the capability of generating continuous words.
We have provided Tab. 3 below, which compares the results of different approaches, namely NeuSpeech, NeuSpeech with teacher forcing, NeuSpeech with pre-training, and joint training. We observe that using NeuSpeech alone often result in semantically coherent sentences, even if not all the words were correct. Particularly in example (3), there was even a hint of expressive tone. However, when NeuSpeech was generating text with teacher forcing, although there were more correct word matches, there were often consecutive word repetitions, such as ”thethe.” The reason for this occurrence is not yet clear, but it significantly affects the semantic coherence of the sentences.
The performance of NeuSpeech improves to some extent when pre-training is employed. This improvement may be attributed, in part, to the model becoming more cautious and refraining from generating further output after a few errors. Among all the examples, only the sentences generated by the pre-trained model had word sequences without any incorrect words. This has both advantages and disadvantages. The advantage is that it improves the precision of the model and reduces the occurrence of incorrect words. However, the disadvantage is that once an error is predicted, the model becomes ineffective.
In contrast, the joint training model seems to tolerate longer sequences of word errors. For example, in (1), the model does not predict a single correct word for most of the sentence but accurately predicts the last seven words, which is neglected by the higher-scoring pre-trained model. This indicates that, even for a model that reaches a certain performance level, we cannot solely rely on scoring metrics. Each model may have unique attributes that we can exploit. In addition, previous approaches for EEG-to-text translation seems to be less effective, as even with the use of teacher forcing, they produced only a few words that match the true labels. This further highlights the superiority of our approach.
In summary, the analysis of generated samples from different approaches reveals valuable insights. NeuSpeech demonstrates the ability to generate semantically coherent sentences, while NeuSpeech with teacher forcing and pre-training offers improvements in word accuracy, albeit with some drawbacks. The joint training model shows resilience to long sequences of word errors. These findings emphasize the importance of considering both quantitative metrics and understanding the unique properties and strengths of different models. Furthermore, the comparison with previous EEG-to-text (Wang & Ji, 2022) methods underscores the advancements achieved by our approach.
4.5 Ablation Study
To demonstrate the effectiveness of our model and provide experimental and theoretical foundations for future improvements in model design, training methods, data augmentation techniques, and modifications, we conducted an ablation study. This study provided a comprehensive analysis of various factors influencing the performance of our model. The experimental results and theoretical insights obtained from these experiments not only validated the effectiveness of our model but also laid the groundwork for designing improved models, selecting appropriate training methods, devising better data augmentation techniques, and making informed model modifications.
4.5.1 Different Dataset
We evaluated our model in Schoffelen dataset (Schoffelen et al., 2019) to illustrate our effectiveness. As shown in Tab. 4, our model has comparable performance on two datasets. This illustrates that our model is applicable under many scenarios, including English and Dutch, and can be adapted on different layouts.
Table 4: Different datasets. Bold for higher socres.
Dataset BLEU-N (%) $\uparrow$ ROUGE-1 (%) $\uparrow$
N=1 N=2 N=3 N=4 F P R
GWilliams 60.3 55.26 51.24 47.78 58.73 60.88 59.76
Schoffelen 52.89 48.53 46.02 44.18 54.28 53.87 56.14
4.5.2 Splits
We conducted further experiments using the GWilliams (Gwilliams et al., 2023) dataset to evaluate NeuSpeech’s generalizability. To test the model’s ability to generalize across recording sessions, subjects, and unique sentences, we developed three additional training/validation splits beyond the standard 8:1 split. Specifically, we withheld: 1) all data from the second recording session, 2) all instances of the story “lw1”, and 3) a random 10% sample of unique sentences. For the latter two splits, we ensured no overlap sentences between the training and test sets for the held-out data.
Table 5: Dataset splits.Noise means input model with Gaussian noise. Random means randomly choosing sentences from training set as predictions for testing set. Retrieve means find best match in training set for each test sentence as predictions using sentence similarity using “all-MiniLM-L6-v2” model. Generation means our model’s generated text. Here, we only shows the BLEU-1 score. Note, for hold-out story and sentences, we used nucleus sampling p=0.25 for generation.
Hold-out Noise Random Retrieve Generation
Session 0.0 4.94 100 53.16
Story 0.0 4.97 11.01 6.91
Sentences 0.0 4.79 16.38 6.54
Table 6: Transcription results on different splits, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings.
Hold-out sentences
Ground Truth: I looked down and saw he was pointing to another marker just slightly different from the one we had just left
Prediction: It was different from now and it just wasn’t working well.
Hold-out story
Ground Truth: Mahogany black curly hair cut short in the typical military style framed a dark skinned face
Prediction: I must admit, this is a great black or cold weather. It had things such as
As shown in Tab. 5, the model maintained good performance when trained and tested on separate recording sessions, demonstrating its ability to generalize across sessions. NeuSpeech significantly outperformed random baselines on splits with no training-test overlap, validating its generalizability despite low text similarity between train and test sentences.Tab. 6 presents sample outputs, revealing the model can still capture some semantic meaning and words when generalizing, though it struggles to produce fully correct novel sentences.
4.5.3 Pretrain
To investigate the effectiveness of pretraining, we conducted experiments involving pretraining on one dataset and fine-tuning on the other. Specifically, we pretrained the model on the Schoffelen dataset (Schoffelen et al., 2019) and fine-tuned on GWilliams (Gwilliams et al., 2023), and vice versa. As shown in Table 7, pretraining provided benefits in some settings.
Table 7: The performance of pretraining. The column Pretrain Dataset means trained on GWilliams, Schoffelen and none. Bold for highest socres.
[…]
N=1 N=2 N=3 N=4
GWilliams Schoffelen 48.09 42.23 38.5 35.60
Schoffelen GWilliams 63.68 61.06 59.26 57.76
- Schoffelen 52.89 48.53 46.02 44.18
- GWilliams 60.30 55.26 51.24 47.78
Interestingly, fine-tuning on GWilliams after pretraining achieved higher performance than training from scratch, while pretraining on GWilliams and fine-tuning on Schoffelen yielded much lower scores. One potential reason for this asymmetry is that GWilliams contains over twice as many samples as Schoffelen. Thus, after pretraining on the larger GWilliams set, the model is already highly optimized, making it difficult for the training loss to further descend during fine-tuning on the smaller Schoffelen data.
In contrast, fine-tuning the model pretrained on Schoffelen Still allowed for improvements to be made on the larger GWilliams distributions. This suggests pretraining can help in low-data regimes but may inhibit learning when moving to a dataset with a broader sample scope. Overall, these pretraining experiments provide insights into how architectural inductive biases interact with dataset characteristics to influence translation performance.
4.5.4 Joint Training
The motivation behind conducting this experiment was to explore whether the simplest approach of padding channels could enable the model to learn shared features from both datasets which is multi-language, multi layout and multi subject. It is important to note that the GWilliams dataset comprises only 208 channels, thus we applied padding on the channel dimensions to match the Schoffelen’s channel dimension of 273. We are concerned that this approach might introduce significant interference, rendering both datasets untrainable. However, our experiment revealed that our worries were unfounded. Despite the decrease in model performance caused by the simple channel padding method, it is indeed effective in integrating the two datasets.
Table 8: Performance of joint training on two datasets.
Dataset BLEU-N (%) ROUGE-1 (%)
N=1 N=2 N=3 N=4 F P R
GWilliams 55.13 52.91 51.28 49.89 57.33 64.22 57.17
Schoffelen 45.12 39.00 35.48 33.03 45.15 46.47 45.49
We got 55.13 and 45.12 BLEU-1 scores on GWilliams and Schoffelen respectively, which is 5-7 points lower compared to training solely from scratch. The results demonstrate that, although the performance was slightly affected, the simple data padding method was able to effectively fuse the two datasets.
However, we acknowledge that the current method we employed is rudimentary, and future work will involve exploring more advanced techniques for dataset integration. This includes incorporating physical information from the sensors and employing more sophisticated approaches to merge the datasets. By doing so, we can further enhance the performance and effectiveness of joint training in multi-language and multi-layout neural transcribing systems.
4.5.5 Scaling Law
To investigate the presence of a scaling law phenomenon in our model, where larger model sizes correspond to improved performance, we conduct an experiment using the Whisper series models at various sizes: tiny, base, small, medium, and large. We limit the GPU resources for training to a single H800 card with 80 GiB of memory, which results in batch size values of 128, 64, 16, 8, and 4, from smallest to largest model.
Figure 1: Performance on different model sizes. Black numbers represent the BLEU-1 score, green numbers are effective epochs after which the evaluation loss does not descend. Note that each experiment runs 120 epochs.
This experiment utilizes the Schoffelen dataset and sets a maximum training epoch of 120. The best-performing checkpoint is selected for each model size. The results, depicted in Fig. 1, indicate that when faced with limited data, increasing the model size leads to a consistent improvement in model performance until a certain size is reached.
This finding suggests that scaling up the model size can be a valuable strategy for enhancing the capabilities and effectiveness of the model in this particular task. However, it is important to consider that larger models may suffer from degradation due to the limited availability of data. Therefore, it is crucial to carefully select an appropriate model size.
Additionally, we observe that the number of epochs required to achieve convergence decreases by approximately half as the model size increases. This observation implies that, even under GPU constraints, increasing the size of the model significantly improves performance without imposing a prohibitive training time penalty. Larger models are able to reach convergence more quickly, despite their increased capacity, resulting in training durations comparable to smaller models. By carefully balancing the bias-variance tradeoff for each specific task and dataset, researchers can fine-tune the model to generalize optimally within the limitations of the available data, thereby maximizing decoding quality given the computational resources at hand.
4.5.6 Data augmentation
There are a lot of research works that used data augmentation in EEG/MEG signal, Bai et al. (Bai et al., 2023) used time-wise masking. Now, we explore three kinds of augmentation, which are masking, noise injecting, and shifting, and compare them with baseline using Whisper-Base in Sec. 4.5.5 which does not have data augmentation. We set three types of masking, which are time-wise, channel-wise, and block-wise. The mask unit is [channel,40],[1,time],[1,40], for the three masking methods respectively. We set two signal-to-noise ratios (SNR), 0dB and 15dB, of Gaussian noise with a probability of 0.5 and 1, which compose 4 settings. We also set 0.5 and 1 probability of data shifting to see if moving the neural data uniformly in the range of max input duration, along with the timestamps can help the model reduce over-fitting. Fig. 2 shows that block masking can improve performance a lot when the mask ratio is relatively low. However, when the masking ratio is bigger, the performance seems to decrease quickly. It can also be noticed that time masking is much worse than the other two mask types to our surprise. It may indicate that the redundancy of the MEG signal is much lower at this sampling rate compared to images which usually use a mask ratio of 75% (He et al., 2022), when the mask ratio is too big to lose essential information, it can be hard for the model to recover information back correctly. Shifting gains no improvements under our experimental setting, oppositely it degrades model performance significantly. To our surprise, adding strong noise which is 0dB can enhance the model’s performance greatly. But for weak noise of SNR 15dB, there is no benefit in our experimental setting. It reminds us that the MEG signal is very noisy, thus using strong noise may let the model better eliminate noise. Therefore, we recommend utilizing block masking with a small mask ratio and low signal-to-noise ratio (SNR) noise for data augmentation as they have demonstrated their effectiveness in enhancing the performance of the models..
Figure 2: Performance changes with different data augmentations. The probability means the likelihood of adding augmentation on each data segment.
4.5.7 Modification
In line with (Défossez et al., 2023), we modified the whisper encoder by replacing the first convolution layer with two separate convolution layers to increase dimensionality and reduce temporal width. Comparing our base model with and without the point-wise convolution layer (Table 9), we observed a significant drop in performance without the layer. This highlights the crucial role of the point-wise convolution layer in achieving good results by capturing important data features and representations.
Table 9: Performance of dropping the point-wise convolution layer. - indicates deletion of module.
Model BLEU-N (%) ROUGE-1 (%)
N=1 N=2 N=3 N=4 F P R
baseline 41.65 36.47 33.51 31.39 44.86 46.71 44.74
- conv 34.81 28.11 24.42 22.01 37.28 39.91 37.17
4.5.8 Data Ratio
We explore the ratio and data used in training, locking validation set and test set to ensure fairness. In Fig. 3, we can see that performance rises as data is adding as expected. Higher performance may come if more data is provided since the curve seems not have reached plateau yet.
Figure 3: Fine-tune layers and data ratio.
4.5.9 Layers in fine-tuning
Also in Fig. 3, it can be observed that the more layers we fine-tune, the better the model will be. Note, that it is only on this base model, and may not be applicable to all.
5 Limitations
While NeuSpeech shows promising results on benchmark MEG-to-text tasks, there remain opportunities to strengthen the scientific validity and real-world applicability of the framework. Future work could focus on incorporating less structured brain signals to better validate translatability of direct human thought, including physical sensor position data for improved accuracy, and exploring augmented datasets to address data scarcity issues that currently limit the model’s ability to generalize and generate completely correct novel sentences. Addressing these limitations through collecting additional training data and enriching the inputs could help advance this emerging field by propelling NeuSpeech towards more precisely decoding thought from brain signals in varied real-world scenarios.
6 Conclusion
This paper explores MEG-to-text using existing speech recognition training, introducing the concept of decoding neural signal as speech. This approach brings high performance on two datasets on standard metrics, and stretches to decode MEG signals without using teacher forcing. We also demonstrate the feasibility of pre-training or jointly training large neural speech models across layouts and languages. We have also explored the effects of augmentation and so on. Despite these advances, the results of hold-out scenarios are relatively low due to the limited number of sentences and the noisy nature of the MEG signal. In our future work, we will solve the problem of limited training sets and also find more mechanisms to fuse multi-modalities, multi-layers with physical information embeddings.
7 Impact Statements
While promising for advancing AI, decoding thoughts from neurodata raises serious privacy, ethical and regulatory concerns if misapplied. We conduct this research openly and transparently, stressing that voluntary informed consent is vital for any future applications. Continued responsible development with strong safeguards is needed to ensure this capability is never abused or used to infringe on individuals’ psychological autonomy, privacy or security without permission. As an emerging field at the interface of neuroscience and technology, clear guidelines and oversight also require discussion to adequately address the risks as this work progresses.